OpenStudy (anonymous):

A baseball player has a lifetime batting average of 0.283. If, in a season, this player has 225 "at bats", what is the probability he gets 73 or more hits? Probability of 73 or more hits

OpenStudy (kropot72):

This can be solved by using the normal approximation to the binomial distribution, where \[mean=np=225\times0.283=63.675\] \[standard\ deviation=\sqrt{np(1-p)}=\sqrt{63.675\times0.717}=6.7568\]